# ADF22+: a declining faint end in the far-infrared luminosity function in the SSA22 protocluster at z=3.09

### v1

@@ 5 Summary
In this paper, we present a deep ALMA 1.17 mm continuum mosaic observation to study the faint DSFG population in an 8 arcmin2 core region of the SSA22 protocluster. The continuum image has resolution and an RMS of Jy beam-1, of the depth in the previous study Umehata et al. (2017). Adopting an S/N cut, we detect 53 continuum sources, doubling the number of detection in the same field. We search for counterparts at other wavelengths using optical-to-mid infrared imaging data of the SSA22 field. For 43 of 44 sources with robust counterparts in NIR bands, redshift, and SFR are also derived from the photometric data. We identify 18 out of 53 sources as SSA22 protocluster members, with spectroscopic redshifts within from .
We compute 1.17 number count and find (1) an excess of 1.17 mm continuum sources at brighter fluxes ( blank fields at mJy), which is produced by the SSA22 protocluster; (2) agreement with blank field surveys after excluding SSA22 protocluster member galaxies. The observed excess of bright 1.17 mm sources is driven by overabundance of massive dusty star-forming galaxies in the SSA22 protocluster core, but SFR in individual member galaxies is not significantly elevated compared to the expectation for their stellar masses, suggesting the SSA22 protocluster core environment enhance the formation of massive galaxies mainly by huge amount of gas supply from the cosmic web instead of frequent major mergers. We also compute the rest-frame 250 m luminosity function of the protocluster core region and find evidence of a declining faint end, mimicking the shape of the bright end optical luminosity function of low-redshift galaxy clusters. These results suggest ongoing build-up of a future galaxy cluster, adding to the growing body of evidence that massive early-type galaxies in local clusters form the bulk of their stellar mass in dusty star formation episodes at high redshift.
@@ Acknowledgements.
This work is supported by the NAOJ ALMA Scientific Research Grant Code 2024-26A and JSPS grant-in-aid No. 22H04939, 23K20035, 24H00004. IRS acknowledges support from STFC (ST/X001075/1). This paper makes use of the following ALMA data: ADS/JAO.ALMA #2013.1.00162.S, #2017.1.01332.S, #2022.1.00680.S and #2023.1.01206.S. ALMA is a partnership of ESO (representing its member states), NSF (USA), and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ.References
@@ Appendix A Source catalogue
Appendix B New spectroscopic redshifts
In this section, we give a brief overview of new spectroscopic redshift measurements. For these sources, the VLT/MUSE spectra are shown in Fig. B.1, B.2 and ALMA band 3 spectra are shown in Fig. B.3. The sources and their spectral features are described below.
@@ Appendix B New spectroscopic redshifts
A48: We find three emission lines for this source: C ii, C iv and C iii] at . The C iv emission line seems broad with an FWHM of km s-1, suggesting AGN activity.
@@
A20: The source has one emission line on the edge of ALMA Band 3 spectrum. The single-peaked photometric redshift of leads to a unique solution as CO(2–1) at .

### v2

@@ 5 Summary
In this paper, we present a deep ALMA 1.17 mm continuum mosaic observation to study the faint DSFG population in an 8 arcmin2 core region of the SSA22 protocluster. The continuum image has resolution and an RMS of Jy beam-1, of the depth in the previous study Umehata et al. (2017). Adopting an S/N cut, we detect 53 continuum sources, doubling the number of detections in the same field. We search for counterparts at other wavelengths using optical-to-mid infrared imaging data of the SSA22 field. For 43 of 44 sources with robust counterparts in NIR bands, redshift, and SFR are also derived from the photometric data. We identify 18 out of 53 sources as SSA22 protocluster members, with spectroscopic redshifts within from .
We compute 1.17 number count and find (1) an excess of 1.17 mm continuum sources at brighter fluxes ( blank fields at mJy), which is produced by the SSA22 protocluster; (2) agreement with blank field surveys after excluding SSA22 protocluster member galaxies. The observed excess of bright 1.17 mm sources is driven by overabundance of massive dusty star-forming galaxies in the SSA22 protocluster core, but SFR in individual member galaxies is not significantly elevated compared to the expectation for their stellar masses, suggesting the SSA22 protocluster core environment enhance the formation of massive galaxies mainly by huge amount of gas supply from the cosmic web instead of frequent major mergers. We also compute the rest-frame 250 m luminosity function of the protocluster core region and find evidence of a declining faint end, mimicking the shape of the bright end optical luminosity function of low-redshift galaxy clusters. These results suggest the ongoing build-up of a future galaxy cluster, adding to the growing body of evidence that massive early-type galaxies in local clusters form the bulk of their stellar mass in dusty star formation episodes at high redshift.
@@ Acknowledgements.
This work is supported by the NAOJ ALMA Scientific Research Grant Code 2024-26A and JSPS grant-in-aid No. 22H04939, 23K20035, 24H00004. IRS acknowledges support from STFC (ST/X001075/1). This paper makes use of the following ALMA data: ADS/JAO.ALMA #2013.1.00162.S, #2017.1.01332.S, #2022.1.00680.S and #2023.1.01206.S. ALMA is a partnership of ESO (representing its member states), NSF (USA), and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. Some of the data presented in this paper were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/d5ma-0904. STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5–26555. Support to MAST for these data is provided by the NASA Office of Space Science via grant NAG5–7584 and by other grants and contracts.References
@@
- Koprowski et al. (2017) Koprowski, M. P., Dunlop, J. S., Michałowski, M. J., et al. 2017, MNRAS, 471, 4155
@@ Appendix A Source catalogue
Appendix B Note on new spectroscopic redshifts
In this section, we give a brief overview of new spectroscopic redshift measurements. For these sources, the VLT/MUSE spectra are shown in Fig. B.1, B.3 and ALMA band 3 spectra are shown in Fig. B.4. The sources and their spectral features are described below.
@@ Appendix B Note on new spectroscopic redshifts
A48: We find three emission lines for this source: C ii, C iv and C iii] at . The source hosts an AGN, as suggested by the broad C iv emission line (FWHM km s-1), detection in the Chandra X-ray data (Lehmer et al. 2009a, b), and a compact morphology in the JWST/NIRCam F444W image (Fig B.2).
@@
A20: The source has one emission line on the edge of the ALMA Band 3 spectrum. The single-peaked photometric redshift of leads to a unique solution as CO(2–1) at .